### v1

# Class Agnostic Instance-level Descriptor for Visual Instance Search

Abstract
Despite the great success of the deep features in content-based image retrieval, the visual instance search remains challenging due to the lack of effective instance level feature representation. Supervised or weakly supervised object detection methods are not among the options due to their poor performance on the unknown object categories. In this paper, based on the feature set output from self-supervised ViT, the instance level region discovery is modeled as detecting the compact feature subsets in a hierarchical fashion. The hierarchical decomposition results in a hierarchy of feature subsets. The non-leaf nodes and leaf nodes on the hierarchy correspond to the various instance regions in an image of different semantic scales. The hierarchical decomposition well addresses the problem of object embedding and occlusions, which are widely observed in the real scenarios. The features derived from the nodes on the hierarchy make up a comprehensive representation for the latent instances in the image. Our instance-level descriptor remains effective on both the known and unknown object categories. Empirical studies on three instance search benchmarks show that it outperforms state-of-the-art methods considerably.
Index Terms:
Instance search, Content-based Image Retrieval, Image search.
Figure 1: Semantic region decomposition in the proposed Class Agnostic Instance-level Descriptor (CLAID). It is a multi-level and multi-granularity instance decomposition, leading to excellent performance on three instance search benchmarks.
I Introduction
Visual instance search is a more challenging search task compared to the conventional content-based image retrieval. It requires the search system to return not only the relevant images, in which the target instance is in presence, but also the localization of the target instance in the relevant images. The localization is usually given as the bounding boxes over the target instance regions. Since the instance only appears as an object in an image sub-region, it could be occluded or undergone various types of transformations. For this reason, features that only describe an image local such as SIFT [1], SURF [2], and DELF [3] are expected. However, although these features are able to represent an object local, a single local feature is usually too fine-grained to form a complete description about an object instance or meaningful part of it (e.g., a logo on a hat). Moreover, there is no way to integrate several local features to achieve that given no prior is provided. Due to the above reasons, the instance-level feature representation is preferred, which is expected to build feature description for a complete instance or a semantic part of an instance.
However, it is non-trivial to design an instance-level feature that is suitable for the search task. First of all, the feature detector, which is responsible for the instance region localization, should be class-agnostic. Otherwise, the detector only performs well on the known object categories [4, 5]. For this reason, supervised or weakly supervised object classification/detection/segmentation models are out of consideration [6]. Moreover, the precise localization on the meaningful instance region or its sub-region is expected. On the one hand, it ensures the bounding boxes returned to the query are in high quality. Moreover, it also guarantees that the instance feature is derived from a precise instance region, which in turn impacts the search quality.
Due to the aforementioned challenges, effective instance-level feature is still slow to occur. In [7], a cycle self-training framework is proposed for instance search, in which the instance search is addressed by an object tracking framework. The instance region is localized by the correlation operation between the query instance and the candidate image as the GlobalTrack [8]. Although good performance is achieved on several benchmarks, such kind of framework is hardly scalable to large-scale search tasks, where the scale of candidate images could be up to million or even billion level. In contrast to self-training framework, DASR (Deeply Activated Salient Region) [4] is able to extract semantically significant regions from the candidate images. The salient regions are obtained by back-propagating peaks on the class activation map of an image. Encouraging search performance is reported on two benchmarks. Unfortunately, the instance localization accuracy turns out to be low in particular for the unknown object categories as the localization in DASR is built upon a pre-trained classification ConvNet.
Recently, the research in computer vision has been marked by the self-supervised ViT model [9]. The distinctive token-wise feature can be built after the self-distillation. These features are so semantically meaningful that they could emerge as the semantic segmentation on an input image. Since the ViT is trained without any label, it becomes the cornerstone for the unsupervised object/semantic segmentation. With the support of self-supervised ViT, each token on the output layer is modeled as a node in an affinity graph, the unsupervised segmentation task is therefore treated as a graph partition problem in [10, 11, 12]. In particular, TokenCut in [12] treats the segmentation as a bipartite graph partition problem. One sub-graph is treated as the foreground object and another is treated as the background. While CutLER in [10] cuts the graph in a chain. Without referring to any other images or labels, significantly better segmentation accuracy than the previous methods is achieved by CutLER and TokenCut.
In this paper, the instance-level regions are detected by a hierarchical decomposition over patch-level features from self-supervised ViT. The decomposition is carried out by repeated bisecting clustering as shown in Figure 1. This process essentially segments one image repeatedly into semantically salient regions. Different from CutLER and TokenCut, we do not simply treat a segmented sub-region as background or a foreground object. As long as they are sufficiently compact, the segmented regions are good candidates of instance-level descriptors. For this reason, the repeated bisecting decomposition continues on the resulting sub-regions until the carefully designed termination condition is reached. The hierarchical decomposition well addresses the problem of object occlusion and object embedding, which are widely observed in the real scenarios.
In summary, the contributions of this paper are threefold. 1) we model the problem of instance-level feature detection as a hierarchical decomposition on an image. The decomposition is addressed by stochastic optimization. This leads to an efficient and relatively exhaustive description about the instances/sub-instances latent in an image. 2) “dummy node” and “terminate condition” are proposed to support the design of this instance-level feature detector. This detector is able to detect instances in different granularity and it is class-agnostic. 3) It shows superior performance on single instance search, multi-instance search as well as image search. This does indicate our method allows the image search and instance search to be unified under the same retrieval framework. To the best of our knowledge, this is the first feature could achieve so in the literature. We have made the codes of our work publicly available [^1] https://github.com/wlzhao22/CLAID.
II Related Work
II-A Instance Search
Before the emergence of deep ConvNet, instance search is addressed as a sub-image retrieval task [13]. Image local features such as SIFT [1] and SURF [2] were widely adopted. Essentially, the search seeks for the correspondence of local features between the query instance and the candidate images. To speed-up the match, feature aggregation methods such as BoVW [14] and VLAD [15, 16] are introduced. Although much better efficiency is achieved, the matching of local features has been degraded to the measurement of similarity between the query instance and the whole candidate image. In addition, due to the sensitivity of local features to the object deformation and 3D rotation, such kind of search paradigm only works on rigid objects.
Due to the wide success of neural-networks in many multimedia tasks, deep features have been introduced to the instance search task [17, 18, 19, 20, 21, 22, 23] in the recent decade. Usually, an image is represented by a deep global feature, which is pooled from certain conv-layer. During the pooling, the higher weights are assigned to the potential instance regions. The feature turns out to be robust to various image transformations, whereas, lacks of distinctiveness. In addition, methods of this category are unable to localize the instance from the relevant images.
In the recent studies [24, 7], the similarity between the query instance and the candidate images is learned via a ConvNet. The correlation is performed between the query instance and each candidate image. High similarity is returned when the target instance appears in the image. Superior performance is reported on several benchmarks. However, the instance search is essentially performed via a deep convolution between the query and all the candidate images. It is hardly scalable to large-scale due to the high time costs.
Instead of borrowing features designed for other tasks, several attempts [25, 4] have been made to design instance-level feature for the search task. In general, there are two major steps involved in the methods, namely instance localization and feature extraction from the detected instance regions. Method in [25] is built upon Faster-RCNN [26], which limits its capability in detecting object from unknown categories. In contrast, DASR [4] relies only on a pre-trained image classification network. It shows much stable performance over [25] as it remains effective on the unknown object categories. However, the localization is imprecise as the detection fully relies on the class activation map. Although encouraging performance is achieved for method in [27], its performance is still unsatisfactory as its region detector relies on weakly supervised learning.
II-B Unsupervised Semantic/Object Segmentation
Owing to the success of self-distilled ViT DINO [9], the high quality of unsupervised semantic/object segmentation/detection is achievable in the recent studies [10, 11, 12]. In these works, the segmentation task is fulfilled by decomposing an affinity matrix, which is built based on the patch-level features derived from DINO. In [11], spectral clustering is adopted for the decomposition. The major eigenvectors after the decomposition correspond to the foreground object in the image. In TokenCut [12] and CutLER [10], the affinity matrix is transformed into a graph by filtering out edges with low similarity. Thereafter, NCut [28] is adopted for the graph partition in both methods. The difference is that only one cut is applied on the graph in TokenCut. The image region that corresponds to the sub-graph with high energy is viewed as an object. While the other is treated as the background. CutLER applies TokenCut three times. The resulting background of each round is cut into foreground object and the background in the next round. Compared to TokenCut, considerably higher segmentation accuracy is achieved. The coarse segmentation formed by the decomposition is further refined by pairwise CRF [11] or MaskCut [10].
Intuitively, methods in unsupervised semantic/object segmentation can be directly borrowed as the detector for the instance search. However, these two tasks are still essentially different. First of all, one cannot assume that the submitted instance query is a complete object. The visual query could be a meaningful sub-region of an object, i.e. a logo on the coat or a cap of the bottle. Or on another extreme, the query is comprised by several visual objects, which is known as multiple-instance search. Theoretically speaking, any semantic part of an image could be cropped out as the query. The instance level feature detector should detect out these semantic regions from one image as much as possible. Different from TokenCut [12] and CutLER [10], a novel hierarchical decomposition procedure is designed in our paper. Given an image is represented by a collection of patch features output from a ViT layer, one feature corresponds to a patch on the ViT input layer. A bisecting procedure is proposed to partition the feature set repeatedly into a bunch of subsets. Each subset corresponds to a semantic region in the image. Instead of simply fixing the number of layers of such bisecting, a condition is designed to judge whether the bisecting should terminate at certain subset.
As a result, similar as the conventional local feature, the collection of such features makes up a full description for the whole image. Whereas, unlike the conventional local feature, the descriptions are instance-level and semantically meaningful. More importantly, in contrast to conventional local features, typically only around 30 descriptors are extracted from one image, which largely relieves the burden of afterwards indexing and nearest-neighbor search.
III Instance-Level Feature Detector
III-A Feature Detection by Hierarchical Decomposition
In ViT [29], an image is partitioned into small patches in equal size like a jigsaw puzzle, an instance inside the image is composed by dozens of such small patches. When feeding into the backbone, one patch is represented into a feature vector. Given $X$ is the set of features produced by the backbone, an instance inside the image corresponds to a subset of features. According to  [9], the pairwise feature distances within the subset are expected to be small, while their distances to the features outside the subset should be relatively larger. Therefore, detecting semantically salient regions could be modeled as a problem of clustering that minimizes the intra-cluster distances. Given there are k clusters $\{S_{1},\cdots,S_{k}\}$, our objective is
[…]
Although the objective looks similar to that of traditional k-means for the first sight, it is essentially different from traditional k-means and most of its variants. The objective minimizes the summation of pairwise $l_{2}$-distance within each cluster. Intuitively, we have to try out all the possible combinations of the feature vectors across different clusters to seek for the optimal solution for Eqn. 1, which is however computationally prohibitive. Moreover, the egg-chicken loop in the traditional k-means is no longer applicable as no cluster centroid is defined in Eqn. 1.
In our paper, the procedure proposed in k-sums [30] is employed to address this optimization problem. Given $x_{i}\in S_{w}$, the overall distance between sample $x_{i}$ and cluster $S_{w}$ is defined as
$d(x_{i},S_{w})=\sum_{x_{j}\in S_{w}}\parallel x_{i}-x_{j}\parallel^{2},$ (2)
which is the summation of distances between feature $x_{i}$ and the rest of features in $S_{w}$. Eqn. 2 can be further simplified as
$d(x_{i},S_{w})=n_{w}{\cdot}x_{i}^{\prime}{\cdot}x_{i}-2{\cdot}x_{i}^{\prime}{\cdot}D_{w}+\sum_{x_{j}\in S_{w}}{x_{j}^{\prime}{\cdot}x_{j}},$ (3)
where $D_{w}=\sum_{x_{j}\in S_{w}}x_{j}$ and $n_{w}=|S_{w}|$. The first term and the last term in Eqn. 3 can be calculated conveniently given the $\textit{l}_{2}$-norms of each feature is pre-computed. The second term is the inner-product between feature $x_{i}$ and $D_{w}$, which is comparable to measuring the distance between $x_{i}$ and a centroid in the k-means. Since features are $l_{2}$-normalized in our problem, Eqn. 3 is simplified further as $d(x_{i},S_{w})=2{\cdot}n_{w}-2{\cdot}x_{i}^{\prime}{\cdot}D_{w}$. Given currently $x_{i}\in S_{w}$, we want to check whether Eqn. 1 would decrease assuming $x_{i}$ was moved to $S_{v}$ already,
$d(x_{i},S_{v})=(n_{v}$ $+1){\cdot}x_{i}^{\prime}{\cdot}x_{i}-2{\cdot}x_{i}^{\prime}{\cdot}(D_{v}+x_{i})$ (4)
$+\sum_{x_{j}\in S_{v}\&j\neq i}{x_{j}^{\prime}{\cdot}x_{j}}+x_{i}^{\prime}{\cdot}x_{i}$
$=n_{v}{\cdot}$ $x_{i}^{\prime}{\cdot}x_{i}-2{\cdot}x_{i}^{\prime}{\cdot}D_{v}+\sum_{x_{j}\in S_{v}\&j\neq i}{x_{j}^{\prime}{\cdot}x_{j}}.$
If $d(x_{i},S_{v})<d(x_{i},S_{w})$, such movement is favored. In particular, we can check over $k-1$ clusters to seek for the most greedy movement that Eqn. 1 decreases the most. In one round, each feature $x_{i}$ is selected by random to seek for the most greedy movement. While if $d(x_{i},S_{v})\geq d(x_{i},S_{w})$, the movement will not be undertaken. The greedy optimization usually converges to a local optima [30]. Since the step that seeks the best movement is comparable to seeking the closest centroid in traditional k-means, the computational complexity of above optimization is on the same par as the traditional k-means.
Repeated Bisecting The primary goal of clustering is to distinguish different instance regions. However, it is not appropriate to segment the image directly into k regions. First of all, the hyper-parameter k is hard to be set. The number of potential instance regions varies from one image to another. Moreover, our perception on visual instances is in multiple scales. For instance, the logo on the bottle is an instance. However, it is only a part of another instance when we view the bottle as an instance. Furthermore, when the bottle is on a desk, the bottle is just a small instance embedded in another bigger instance “desk”. In order to mimic the human perception about the visual instances, the clustering is carried out hierarchically. Specifically, the feature set is bisected into two each time. For a complete image as input, each step we decompose the image or its sub-image into two fine-grained sub-regions. Different levels of decomposition locate visual instances of different scales. A termination condition, which will be detailed later in this section, is designed to decide whether the bisecting should not be carried out further on a subset. By this way, we avoid setting the hyper-parameter k explicitly. The number of clusters we finally have is largely related to the number of latent instances in an image.
Given a node on the hierarchy, one more step is undertaken to partition the resulting subsets further into smaller ones based on the spatial connectivity of patches. Therefore, the hierarchy is not a binary tree in the usual case.
In the following, the details are presented about how the hierarchical detection procedure is initialized and what is the condition to terminate the bisecting on a node. It is possible that the segmented regions are the empty background or the loose structures. A condition is further designed to filter out these dummy nodes on the hierarchy.
III-B Decomposition Seed Selection
Given a feature set (or a node) $S$, the bisecting clustering starts from two randomly bisected clusters on $S$ [30]. Due to the random initialization, the clustering result varies from one run to another. To address this issue, a prior is introduced to initialize the bisecting. To facilitate our discussion, affinity matrix A is defined as follows
$A_{ij}=\begin{cases}1,&x_{i}^{\prime}{\cdot}x_{j}>\alpha,\\ 0,&otherwise,\\ \end{cases}$ (5)
where $x_{i},x_{j}\in S$, and $\alpha$ is empirically set to 0.2[31]. Given $x_{b}\in S$ is the feature holds the highest degree according to A (given in Eqn. 6) and $x_{w}$ (given in Eqn. 7) is the feature holds the lowest degree, they are most likely located on different instances as discovered in [31].
[…]
So $x_{b}$ and $x_{w}$ are selected as two seeds for the bisecting. The rest of features in $S$ are assigned to cluster of $x_{b}$ or $x_{w}$ by checking whether they are closer to $x_{b}$ or $x_{w}$. The initialization scheme is adopted for the bisecting of all the nodes on the hierarchy. The difference between the clustering with seed selection and clustering with random initialization is illustrated in Figure 2.
Figure 2: The process diagram showcases different initialization methods for the bisecting clustering. The top row illustrates the clustering with seed selection. The first column shows the original image, followed by the images after the initialization, intermediate clustering result, and the final clustering results. The bottom row shows the clustering under random initialization.
III-C Terminate Condition
Given a feature set (or a node) $S$ on the hierarchy, the bisecting over it should not be carried out when the connectivity within the feature set is already sufficiently tight. The integral of these features inside usually represents a compact semantic region in the image. Given the total number of edges within the feature set $S$ is
$c(S)=\sum_{i,j\in S\&i<j}A_{ij},$ (8)
the average internal connectivity within $S$ is defined as
$\bar{c}(S)=2\cdot\frac{c(S)}{n\cdot(n-1)},$ (9)
where $n=|S|$. Apparently, $\bar{c}(S)$ is in the range $[0,1]$. During the top-down decomposition, as the bisecting is carried out on a deeper node, the feature $S$ becomes more compact. $\bar{c}(S)$, therefore, gets higher. We can control decomposition granularity by tuning the threshold $\tau_{1}$ on $\bar{c}(S)$. A larger threshold leads to finer decomposition granularity. In our implementation, when $\bar{c}(S)\geq{\tau_{1}}$, the bisecting on $S$ is not undertaken further across all the experiments.
III-D Condition on Dummy Nodes
Given a feature set (or a node) $S$ on the hierarchy, it does not necessarily correspond to a semantically compact region. It could be a mixture of multiple instances of different categories or the mixture of objects and the background, or simply empty background. Such feature set is called “dummy node”. It is not meaningful to build instance-level feature from it. Inspired by [32], “dummy node” is judged by the strength of feature energy. In our method, we define $H=\arg\max\nolimits_{\forall x_{i}\in X}^{\theta}\left|x_{i}\right|^{1}$ as the set of features that exhibit high energy within the dataset $X$. Here, we set $\theta$ to the top-$30\%$ based on the relevant experimental results presented in [32]. We measure the overlapping ratio between $S$ and $H$
$\xi(S)=\frac{\left|H\cap S\right|}{\left|S\right|}.$ (10)
When $\xi(S)$ is low, it basically indicates the region is over-dominated by semantically insignificant features. A threshold $\tau_{2}$ is set to ignore these dummy nodes on the hierarchy. No feature will be built from it when a detected region is viewed as “dummy node”. However, this condition does not prevent the feature set $S$ from further bisecting. It is possible that the direct/indirect subset of $S$ becomes a salient region when the bisecting reaches a more fine-grained level.
Figure 3: The framework of building instance-level feature for instance search. There are two major components, instance-level region detector and feature descriptor. Given an image, a set of patch-level features are produced by the self-supervised backbone. The hierarchical decomposition is applied on the feature set. Each node on the hierarchy corresponds to a potential instance region or sub-region. The patch-level masks can be produced for the valid nodes on the hierarchy. The feature for each region is pooled from another network. Given an incoming query, feature is extracted with the same backbone.
Input: Input image $I$
Output: Set of regions $\mathcal{R}$.
1 Initialize $X=backbone(I)$ ;
2 Initialize $\mathcal{L}\leftarrow X$ and $\mathcal{R}\leftarrow\emptyset$;
3 while $\mathcal{L}\neq\emptyset$ do
4      Pop $node$ from $\mathcal{L}$;
5       if $\xi(node)>\tau_{2}$ and $node\neq root$ then
6            $\mathcal{R}$.append($node$);
7       end if
8      if $\bar{c}(node)<\tau_{1}$ then
9            $S_{b},S_{w}\leftarrow$ k-sum($node$);
10             $\mathcal{L}\leftarrow\mathcal{L}\cup\text{getObject}(S_{b})\cup\text{getObject}(S_{w})$;
11
12       end if
13
14 end while
Algorithm 1 Instance-Level Feature Detector
The entire process of instance-level detection is summarized in Algorithm 1. The hierarchical decomposition starts from feature set $X$. The bisecting is repeatedly applied on $X$ and its subsets. The resulting subsets $S_{b}$ and $S_{w}$ do not necessarily correspond to two instance regions. Several instances from the same object category are likely partitioned into one subset. As a result, one more step “$getObject(\cdot)$” (Algorithm 1 Line – 10) is undertaken to partition subsets $S_{b}$ and $S_{w}$ respectively into spatially connected smaller subsets. Two subsidiary conditions are involved in the algorithm. Specifically, if a node is dummy according to $\xi(node)$, it will not be joined into the region list. $\bar{c}(node)$ determines whether a node should be bisected further.
Discussion The advantages of hierarchical decomposition are at least three folds. Firstly, it well addresses the issue of object embedding and occlusion. In addition, the threshold $\tau_{1}$ is easier to set than the hyper-parameter k. The detected regions are the well reflection of the latent instance distribution in one image. The top-down decomposition reduces the bias accumulation caused by the bisecting and decreases the scale of the partition as the bisecting goes deeper. In the bisecting procedure, the partition could be achieved by NCut [28]. Nevertheless, the bisecting clustering is favored over NCut as the latter is much slower. Moreover, the decomposition by bisecting clustering is favored because it adapts well to the multi-scale semantic structure of an image, which will be revealed in the experiments.
TABLE I: Search performance (mAP) on Instance-160, Instance-335, and INSTRE. digits are collected from the referred paper
Method Descriptor Backbone Instance-160 Instance-335 INSTRE
mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all mAP-50 mAP-100 mAP-all
R-MAC [17] VGG-16 0.302 0.281 0.330 0.417 0.354 0.381 0.741 0.626 0.523
CAM-weight [19] VGG-16 0.337 0.310 0.358 0.404 0.336 0.347 0.521 0.388 0.320
BLCF [20] VGG-16 0.641 0.617 0.651 0.506 0.460 0.483 0.657 0.578 0.636
BLCF-SalGAN [20] VGG-16 0.633 0.604 0.638 0.490 0.442 0.470 0.821 0.751 0.698
DASR* [4] ResNet-50 0.752 0.720 0.771 0.783 0.705 0.724 0.870 0.779 0.692
CutLER [10] ResNet-50 0.675 0.635 0.683 0.712 0.630 0.641 0.870 0.799 0.727
DUODIS [27] VGG-16 - - - - - - - - 0.667
DUODIS [27] ResNet-101 - - - - - - - - 0.752
CLAID (NCut) ResNet-50 0.771 0.732 0.780 0.785 0.703 0.722 0.889 0.822 0.765
CLAID VGG-16 0.686 0.635 0.681 0.674 0.581 0.591 0.851 0.755 0.667
CLAID ResNet-50 0.781 0.743 0.791 0.789 0.710 0.725 0.894 0.827 0.775
CLAID ResNet-101 0.786 0.745 0.797 0.793 0.715 0.735 0.893 0.825 0.773
IV Feature Descriptor
The output from the repeated bisecting on the feature set $X$ is a hierarchy. Each node on the hierarchy is a subset of $X$ (except for the root node), which corresponds to a potential salient regions. In this section, we show how the features are extracted for these detected salient region, except for the dummy nodes on this hierarchy (discussed in Section III-D).
Since the features are on the patch level, there is a one-by-one correspondence between the feature and a patch on the ViT input layer. The segmented subset on the hierarchy appears as a bunch of connected patches on the image. The detected region is given as a region with a patch-level mask. Unlike unsupervised object detection or semantic segmentation [10, 31], fine-grained localization is not necessary for instance search. Therefore, the post-processing steps such as up-sampling and boundary modification of location results, such as CRF [33] are not adopted in our work.
Given the location of the salient region is available, the feature extraction is as easy as performing RoI-pooling on a feature map from a network layer. To this end, we have multiple options. For the network, we can use the self-supervised ViT [9] based on which our salient regions are detected. Alternatively, the pre-trained ConvNet e.g., VGG [34] or ResNet [35] can be used to derive the feature. Moreover, instead of feeding the entire image into the network and pooling based on the region box, the detected region can be cropped out and fed into the network as one image to produce the feature. However, this way is computational inefficient. For this reason, we choose to feed the image into the pre-trained ResNet and extract the instance-level feature via RoI-pooling.
The complete pipeline of our hierarchical region detection and feature extraction is shown in Figure 3. On the first stage, the potential instance regions or sub-regions are detected via a top-down hierarchical decomposition over the feature set produced by backbone. Given the detection is an unsupervised procedure and is built upon class-agnostic self-supervised ViT and the feature is extracted from a ResNet. Since the detector is built upon class-agnostic self-supervised ViT the produced feature becomes a class-agnostic descriptor in the real sense. For this reason, our feature is called CLass Agnostic Instance-level Descriptor (CLAID) from now on. In contrast to traditional local feature, only 30 instance level regions per image are extracted.
V Experiment
V-A Experiment Setup
The effectiveness of the proposed instance-level feature CLAID is evaluated on three benchmarks for instance search. They are Instance-160, Instance-335222https://drive.google.com/file/d/1ZxVUO3uInFqiFzXfuokQxalwMl1fEDgN
/view?usp=drive_link
and INSTRE [36]. Instance-160 contains 160 queries and 11,885 reference images. Instance-335 is an augmented dataset of Instance-160 with data collected from the GOT-10K [37], Youtube BoundingBoxes [38], and LaSOT [39] datasets. There are 335 instance queries and 40,914 reference images. For INSTRE dataset, there are 1,250 queries and 27,293 reference images. Compared to Instance-160 and Instance-335, INSTRE is rich of background variations and covers a wide range of object categories. In particular, there are 250 queries containing multiple instances. In order to verify the performance of CLAID in the multiple instance search task, the performance on this subset of queries (given as INSTRE-M) is further studied in the paper. The search quality is measured by mAP. Specifically, mAP-50, mAP-100, and mAP-all are reported for our method and state-of-the-art methods considered in the comparison. In addition, we also reported the instance localization accuracy on three datasets, which is measured by the commonly used Intersection Over Union (IoU). In order to study the exhaustiveness of our detector in discovering the latent instance regions in one image, we test it as an unsupervised object detection method. We study the recall of the correct localizations on COCO validation dataset [40] in comparison to state-of-the-art methods CutLER [10], TokenCut [12], and Deep Spectral [11].
Theoretically speaking, detector can be built upon any ViT networks. Due to its outstanding performance, ViT-Base (ViT-B/8) trained by DINO [9] is selected as the default backbone. For feature extraction backbone, there could be several options as well. ResNet-50  [35] is selected in our standard configuration follwoing the same setting in  [4] where feature is derived from Block4.0. Moreover, the search performance based on the other backbones such as ResNet-101 [35] and VGG-16 [34] is presented to validate the effectiveness of our detector. Both the detector and descriptor were implemented using PyTorch. The experiments were conducted on an Nvidia RTX 4090 GPU.
Figure 4: This display showcases the results of instance search by CLAID. The instance queries are from images enclosed in a blue border. Each query is followed by four retrieved images from the top-1, top-5, top-10, and top-50 outcomes for that query. True-positives are colored in green, while false-positives are in red. CLAID performs well in the majority of queries. Nevertheless, its performance degrades on blurry images or semantically specific queries (e.g., a specific human face).
TABLE II: Performance (mAP) on multi-instance search task
Methods mAP-50 mAP-100 mAP-all
R-MAC [17] 0.824 0.742 0.465
CAM-weight [41] 0.593 0.464 0.267
BLCF-SalGAN [20] 0.938 0.914 0.750
DASR* [4] 0.962 0.916 0.692
CutLER [10] 0.964 0.942 0.768
CLAID 0.964 0.943 0.774
TABLE III: Localization accuracy (mIoU) on Instance-160, Instance-335, and INSTRE. The IoU threshold is on 0.5 level
Methods Instance-160 Instance-335 INSTRE
DASR [4] 0.252 0.299 0.417
DASR* [4] 0.260 0.286 0.369
CutLER [10] 0.262 0.488 0.709
CLAID 0.377 0.527 0.696
V-B Instance Search Performance
In the first experiment, the effectiveness of our method is studied in comparison to several representative methods in instance search. Among these methods, R-MAC [17], CAM-weight [41], BLCF [20] and BLCF-SalGAN [20] are deep global features which collapse all the features of an image into one vector for retrieval. In contrast, methods like DASR [4] and ours extract the instance level feature from the candidate images and are able to localize the instance from the candidate images. In addition, we also treat the unsupervised object detector CutLER [10] as an instance region detector. The same feature as ours is built for the regions detected by CutLER. Moreover, we also report the results when the bisecting clustering in CLAID is replaced by NCut [28], while keeping other configurations unchanged. The instance search performance is shown in Table I.
As seen from the table, global features show poor performance on the instance search task. Object level features such as DASR, CutLER, and CLAID demonstrate considerably better performance. CLAID outperforms the rest of methods on three benchmarks. Particularly, CLAID shows considerably better performance than the rest on INSTRE where the queries are more challenging and there are a lot of distractors in the reference set. Although CutLER is also able to build instance level features, its performance on instance search is considerably lower than CLAID. The relatively poor performance mainly dues to its limited number of partitions over the feature set. Moreover, the cut in a chain in CutLER also causes the loss of potential instance regions. Compared to CLAID (NCut), CLAID performs considerably better. According to our observation, the results from bisecting clustering better reflect the multi-scale semantic structure than the partitions produced by NCut. In addition, we also tested different backbones with CLAID at its feature extraction stage. As shown from the table, our method shows considerably better performance when it is supported by the same backbone as the competing methods (e.g., DASR and CutLER). This largely owes to its high localization accuracy. Figure 4 further showcases the instance search results from CLAID.
In order to study the behavior of our method on multi-instance search task, the performance on 250 queries from INSTRE is particularly evaluated. For these queries, there are more than one objects in presence in a query. As shown in Table II, most of the methods show relatively better performance on this task than that of single instance search (reported in Table I). The search quality from all global features is considerably better. This is mainly because the large coverage of the instance query over the image makes it close to a conventional content-based image search task. The mixture of features from different instances becomes helpful in this context. Our method still shows the best performance as it keeps the features extracted from different scales. The localization accuracy of different methods (measured by mIoU) is reported on Table III. Since the localization is impossible to fulfill for the global features, only the performance from instance level features is reported. The patch-level region masks from our method and CutLER are transformed to bounding boxes when we calculate the mIoU score. As shown in Table III, CLAID shows considerably better localization accuracy over DASR* and CutLER.
V-C The Performance of Feature Detector
Figure 5: Recall at different IoU thresholds for four methods on COCO 20k.
Since the feature detector is the critical component in our pipeline, its effectiveness is further studied from two perspectives, namely the detection speed and the exhaustiveness in detecting latent instance regions from an image. The processing speed is compared to unsupervised object detectors such as Deep Spectral [11], TokenCut [12], and CutLER [10]. All of them are built on self-supervised ViT. For this reason, we only evaluate the time cost of processing one image. The speed efficiency is evaluated on CPU with single-thread.
The average number of cuts that are carried out on one image, the average time of processing one image, and the time cost of one cut are shown in Table IV. In TokenCut, CutLER, and CLAID (NCut), NCut [28] is adopted to carry out the partition. The NCut is computationally intensive as it involves the eigenvalue decomposition on a matrix in big size. The time cost of one cut is much higher than that of the bisecting clustering adopted by CLAID. The processing time for CLAID (NCut) is relatively shorter than the other methods based on NCut. The cut time becomes shorter since the cut is applied on the small subsets on the hierarchy. Owing to the efficiency of bisecting clustering, CLAID is 6$\times$ faster than CLAID (Ncut). To verify the ability of our detector in detecting the latent instance regions in an arbitrary image, we further study the performance of CLAID on COCO 20k segmentation validation dataset. In this evaluation, CLAID is treated as an unsupervised instance detection method. The exhaustiveness is measured by the recall rate of the object instances in different categories that have been annotated in the dataset. The recall is evaluated with respect to mIoU score. Three detectors Deep Spectral [11], TokenCut [12] and CutLER [10] are considered in the comparison. The detection performance is shown in Figure 5. As seen from the figure, CLAID outperforms these detectors by a large margin because the hierarchical decomposition is able to discover more correct object regions.
TABLE IV: The comparison on the processing time costs
Methods ##\##cut TM per image (s) TM per cut (s)
Deep Spectral [11] 5 17.3 3.457
TokenCut [12] 1 9.6 9.600
CutLER [10] 3 28.7 9.566
CLAID (NCut) 29 14.6 0.503
CLAID 29 2.40 0.083
V-D Performance on Image Retrieval
Generally speaking, image retrieval can be viewed as a multi-instance search task. Our method is, therefore, also feasible for conventional image retrieval. In this experiment, we study the effectiveness of CLAID on image retrieval. Following the same processing pipeline, a bunch of CLAID features are extracted from one candidate image. While only one feature is extracted from a query image. The ranking list is produced based on the position where a candidate image first appears in the instance search ranking list.
Performance is evaluated on three benchmarks Paris6k [42], Holidays [43], and Oxford5k [44]. As shown from the Table V, CLAID consistently outperforms a majority of representative methods in the literature. The best performance is achieved by the DUODIS [27]. DUODIS excels by mining latent objects directly from the test datasets and employs a pre-trained model, which is pre-conditioned on images of landmarks and buildings for feature extraction.
TABLE V: Performance (mAP) on image retrieval. digits are from the referred paper. digits are from  [4]
MEthods Holidays Oxford5k Paris6k
R-MAC [17] - 0.669 0.830
CAM-weight [19] 0.785 0.712 0.805
BLCF [20] 0.854 0.722 0.798
BLCF-SalGAN [20] 0.835 0.746 0.812
DASR* [4] 0.873 0.613 0.744
DUODIS [27] - 0.900 0.952
CLAID 0.899 0.780 0.872
V-E Ablation Studies
To study the impact of each major component in CLAID, an ablation analysis is carried out on INSTRE. Namely, we study the impact of seed selection (Section III-B) and the contribution from the filtering scheme on the “dummy node” (Section III-D). Table VI showcases the mAP results and the total number of features extracted from INSTRE images under each configuration. As shown in the Table, compared to CLAID with random initialization, the detector produces 16% less redundant instance regions. Moreover, compared to the run extracting features for dummy nodes, the standard configuration of CLAID shows slightly higher mAP while using 42% less number of features. This basically indicates that regions that correspond to the dummy nodes make no contribution to the instance level feature representation.
TABLE VI: Ablation study on the seed selection and the filtering strategy on INSTRE
Init. dummy INSTRE
mAP-50 mAP-100 mAP-all #Features
seed $\times$ 0.893 0.826 0.774 1,180,746
rand $\surd$ 0.894 0.827 0.775 962,154
seed $\surd$ 0.894 0.827 0.775 827,426
V-F Sensitivity Studies
In our method, there are two key hyperparameters namely the average internal connectivity ($\tau_{1}$) and the level of semantic dominance ($\tau_{2}$). To evaluate the impact of hyperparameters settings on our model’s performance, a series of sensitivity tests have been pulled out. The experiments are carried out on Instance-160. In the experiments, when validating $\tau_{2}$, we fix $\tau_{1}$ to 0.97. Conversely, when validating $\tau_{1}$, we fixed $\tau_{2}$ to 0.2. This setup allows us to isolate the impact of each parameter and analyze their individual impact on the model’s performance. The instance search performance trends on Instance-160 are presented in Figure 6. As shown from the figure, the search performance is more sensitive to the level of semantic dominance ($\tau_{2}$), while only minor performance variation is observed when $\tau_{1}$ varies. Nevertheless, the search performance remains stable as long as $\tau_{2}$ is in the range $[0.2,~{}0.5]$.
Figure 6: Experimental results of mAP with varying $\tau_{1}$ and $\tau_{2}$ settings on Instance-160.
VI Conclusions
We have presented a novel instance level feature descriptor, namely CLAID. Given an image is represented by the feature set produced by self-supervised ViT, a novel top-down repeated bisecting procedure is proposed to partition the feature set into compact and spatially connected instance regions. The nodes on the resulting hierarchy correspond to salient instance regions mined from the image. Compared to the existing methods, it provides a more exhaustive description about the latent instances in one image. Most importantly, our framework is class-agnostic, which allows it to remain effective on various object categories. Excellent performance is achieved on single and multiple instance search tasks as well as image search task. It, therefore, becomes possible to unify the visual instance search and the content-based image search under one framework. Moreover, only around 30 features are extracted per image, it allows the search to be scaled up to billion images level on a single server.